# uDistil-Whisper: Label-Free Data Filtering for Knowledge Distillation in Low-Data Regimes

### v3

@@ 5 Results and Discussion
$\quad$-- proxy 35.1/17.1 37.2/18.7 58.9/33.6 55.1/31.4 60.3/23.8
$\quad$-- sonar 35.4/16.9 37.7/18.5 57.9/33.1 54.7/31.3 60.9/24.8
Table 2: Average WER/CER across different evaluation datasets. Best results are shown in bold. Second best results are underlined. WER/CER scores after normalization and removing diacritics.
Scaling the Data.
We scale our dataset from 100K to 500K segments and we train two models, UDW-16-16 and UDW-32-16, employing proxy-ref and sonar-sim as filtering criteria. As shown in Table 2, our UDW-32-16++ model, trained with 500K segments, performs comparably with its supervised counterpart DW-32-16++. For instance, UDW-32-16++ with proxy-ref filtering achieves a WER of 51.4%, similar to DW-32-16++’s 50.0%. Notably, our smaller student model (UDW-16-16++) trained on 500K segments results in superior performance compared to models trained under supervised filtering. For example, UDW-16-16++ with sonar-sim gives 53.9% WER compared to 56.2 for DW-16-16++. These results showcase the efficacy of our data filtering methods.
Generalization to Unseen Dialects.
Model Split NJD MTOS KHLJ HJZ UNK
W-FT Test 106.4/77.5 77.0/51.8 117.8/85.4 84.2/61.5 139.9/112.2
Valid 77.1/52.6 63.4/41.1 139.4/100.3 119.1/89.7 140.3/107.6
SM4T-v1 Test 56.5/30.9 74.0/46.0 61.9/32.2 54.6/29.0 69.6/39.4
Valid 51.9/28.1 68.7/44.0 61.7/31.5 54.2/30.9 62.3/35.2
SM4T-v2 Test 52.6/31.1 76.2/53.1 58.2/30.4 53.3/32.0 74.7/45.1
Valid 52.2/30.7 75.8/53.7 65.1/35.3 51.1/30.3 59.8/34.4
W-M Test 88.2/65.8 107.9/79.3 102.8/77.0 86.7/59.7 134.0/122.2
Valid 80.4/56.9 102.8/75.1 89.5/62.9 72.9/52.0 127.7/106.5
W-L-v2 Test 56.6/39.9 74.0/57.4 80.2/54.4 63.4/39.6 98.2/80.7
Valid 60.9/41.4 72.9/55.4 67.7/44.9 64.5/43.6 68.0/67.1
W-L-v3 Test 53.1/31.6 73.7/53.7 70.5/44.1 61.7/38.6 89.0/61.3
Valid 49.3/30.2 65.5/47.7 67.5/39.2 46.5/27.2 67.7/49.2
DW-16-16 Test 58.2/30.8 74.9/47.6 65.1/32.7 58.2/30.7 68.4/39.8
Valid 59.4/31.4 70.6/44.7 66.2/35.2 61.1/32.8 69.9/39.8
DW-32-16 Test 58.7/35.8 74.9/60.1 66.7/38.7 58.7/34.1 70.1/44.5
Valid 58.3/34.8 69.7/54.1 67.5/37.6 62.7/38.2 68.3/40.1
DW-16-16++ Test 58.8/30.9 81.9/50.7 62.1/31.5 57.5/31.0 77.5/46.8
Valid 56.8/29.8 72.0/43.8 62.3/31.8 60.2/33.0 75.2/41.0
DW-32-16++ Test 51.7/28.3 67.8/43.1 57.5/29.4 52.6/28.6 67.5/41.3
Valid 50.3/27.3 61.8/38.3 62.3/34.5 53.7/28.1 66.4/43.0
@@
$\quad$-- DW-16-16 Test 61.6/34.8 90.4/59.7 72.9/41.4 68.5/42.3 88.8/63.0
Valid 64.8/38.9 80.3/53.4 71.4/41.9 65.5/37.7 77.0/54.3
$\quad$-- DW-32-16 Test 58.2/42.8 82.3/63.9 67.3/47.0 60.4/45.9 76.6/63.2
Valid 57.9/35.2 73.7/54.9 68.6/43.3 56.8/36.5 72.3/49.6
@@
$\quad$-- proxy Test 60.2/35.5 75.6/55.6 65.7/38.9 62.1/39.6 75.6/52.0
Valid 59.3/34.0 70.7/50.9 66.5/39.1 61.4/37.1 68.7/41.2
$\quad$-- sonar Test 64.6/35.8 59.9/30.3 84.6/55.7 72.1/38.8 66.6/36.8
Valid 64.8/35.9 67.5/39.3 78.1/52.4 69.9/38.7 65.3/36.7
@@
$\quad$-- proxy Test 53.4/31.1 75.7/54.0 60.0/32.1 51.4/30.8 72.8/46.0
Valid 51.0/29.3 65.9/44.3 58.1/29.1 53.7/28.6 64.9/36.6
$\quad$-- sonar Test 48.8/25.4 52.6/23.6 67.5/45.9 58.3/29.9 48.9/25.5
Valid 49.2/26.0 51.6/26.7 62.5/44.1 58.9/30.3 52.6/29.5
@@
$\quad$-- proxy Test 53.4/29.7 73.9/48.8 61.9/33.8 53.3/29.6 68.6/42.8
Valid 52.7/27.8 66.6/42.0 62.0/34.3 55.7/32.2 67.3/41.7
$\quad$-- sonar Test 53.7/28.4 69.4/43.3 61.4/30.8 53.7/27.5 64.6/37.0
Valid 53.6/27.5 64.4/40.3 62.3/32.4 55.7/30.7 63.7/35.8
UDW-32-16++
@@
Valid 49.0/25.3 60.4/37.4 57.6/30.1 50.7/25.6 60.7/37.4
$\quad$-- sonar Test 50.5/26.3 65.3/40.8 57.3/28.3 49.1/24.8 61.8/34.5
Valid 49.3/25.3 58.8/37.0 58.1/30.2 53.6/29.8 61.1/34.2
Table 3: Results on top five dialects/categories in SADA data. Best results are shown in bold. Second best results are underlined. WER/CER scores after normalization and removing diacritics.
We evaluate our models under various unseen conditions, including five novel dialects that our models have not encountered before. Additionally, we compute category-wise  [^5] By categories we refer to dialectal and non-dialectal labels in the SADA data. error rates for both the SADA test and valid splits. We present results on the top five categories. As detailed in Table 3, our best model UDW-32-16++ consistently outperforms all others when averaged across these novel dialects.
@@
Evaluation Dataset W-L-v2 DW-16-16 DW-32-16 UDW-16-16-proxy UDW-32-16-proxy
IID OpenBible 101.33/44.43 59.14/13.95 58.78/13.75 59.21/13.96 58.86/14.13
CommonVoice17 117.13/60.13 82.91/34.96 69.8/24.77 75.58/29.18 70.43/25.36
ALFAA 217.11/143.19 78.24/28.23 74.4/25.72 76.82/27.2 73.81/26.53
OOD DVoice 214.6/144.57 124.35/74.1 110.19/62.61 110.68/62.4 114.9/69.14
AMMI-LigAikuma 46.71/12.96 60.1/18.02 51.77/14.37 60.44/18.47 52.19/14.42
Fleurs 54.61/14.81 60.85/18.85 51.58/14.8 58.87/18.45 51.82/14.88
Table 4: WER/CER comparison on Swahili. Best results are shown in bold. Second best results are underlined.
To further validate the effectiveness of our approach, we conduct experiments on Swahili, a low-resource language. We collect over 100 hours of labeled speech data from a variety of sources, namely OpenBible Meyer et al. (2022), CommonVoice (Swahili subset) Ardila et al. (2020), ALFAA [^6] https://github.com/besacier/ALFFA_PUBLIC/tree/master/ASR/SWAHILI, DVoice Gauthier et al. (2016), AMMI-LigAikum [^7] https://github.com/besacier/AMMIcourse, and FLEURS (Swahili subset) Conneau et al. (2022).
@@
We compare our distilled models to the teacher model and evaluate the performance of our unsupervised approach. The results show that our unsupervised distillation models perform on par with, or better than the supervised setup. Additionally, our distilled models outperform the teacher model by a significant margin on both IID and OOD datasets, demonstrating the utility of our approach in extremely low-resource settings. Specifically, the UDW-32-16 model achieves a WER/CER of 58.86/14.13% on the IID OpenBible dataset, compared to the teacher model’s 101.33/44.43%. On the OOD dataset FLEURS, UDW-32-16 attains a WER/CER of 51.82/14.88, significantly outperforming the teacher model’s 54.61/14.81. Across various datasets, our distilled models consistently outperform the teacher, with UDW-32-16 showing the best results overall. Table 4 presents the WER and CER scores for the different models and datasets.
@@ 6 Conclusion
In this study, we explore methods for distilling large Whisper models into smaller, more efficient ones. Our proposed techniques for filtering out low-quality examples eliminate the need for labeled data, bridging a gap in prior research and facilitating the creation of compact and effective speech recognition models for settings with limited labeled data. We conduct a comprehensive evaluation of our trained models, comparing them against various strong zero-shot and distilled models. Our results demonstrate that our models can outperform both the teacher model and models distilled with supervised data filtering setups.
We evaluate our models across a diverse range of Arabic varieties and dialects, assessing their generalization and adaptability to linguistic diversity. From large multilingual speech models (Whisper), we distill language-specific models optimized for Arabic ASR. Our evaluation spans three different setups, encompassing a wide array of linguistic varieties (ie: dialects), and reveals that despite their reduced size (25-50%), these distilled models consistently outperform their base counterparts and can compete with state-of-the-art models twice their size—all achieved without the need for labeled data. We also apply our unsupervised distillation approach on Swahili, another low-resource language, which shows a similar performance and proves its efficacy in different setting.
Notably, our model-based unsupervised data filtering methods (proxy and sonar) demonstrate superior robustness across linguistic variations. Moving forward, we aim to explore model-free approaches to further enhance the efficacy of filtering low-quality pseudo labels. Additionally, our future research will extend to include extremely low-resource languages and domains.
@@ 7 Limitations
Our main approach relies heavily on a robust reference model to serve as a proxy for filtering lower-quality pseudo labels. Specifically, we utilize SeamlessM4T-large-v2, a state-of-the-art model with 2.3 billion parameters, to generate proxy references which is then used to filter out low-quality data points. For similarity-based measures, we use SONAR to generate multimodal embeddings from speech and pseudo labels. These embeddings provide contextual similarity which is then utilized to discard low-quality pseudo labels. We use AceGPT (7B), to compute the log-likelihood of the pseudo labels which is leveraged to filter out low-quality examples.
@@
Evaluation. Arabic is a linguistically rich and complex language with over 400 million speakers Abdul-Mageed et al. (2021), resulting in its wide range of varieties and dialects. We evaluate all the models on eleven different datasets representing different varieties, including five novel dialects collected and curated by native speakers and never seen before by any models. However, our varieties do not cover all Arabic-speaking regions. We aim to address this in future work by covering more varieties and dialects.
Distillation Training Data. We distilled four variants of student models using 100K and 500K segments of which approximately 25% are filtered. We see improvement going from 100K ($\approx$100100100100 hours) to 500K ($\approx$500500500500 hours) segments. As Gandhi et al. (2023) shows, going over 1000 hours results in a better model, we aim to study how distillation can be done under a low resource setting which is why we do not scale the data. Additionally, we also keep the WER threshold high (80) so that we remain close to a setting where no labeled data is available (even for filtering). It would be interesting, however, to see how distilled models may perform on unfiltered data in low-resource setting.
@@ 2.1 SADA Dataset
Table 5: SADA stats. S is the number of segments and D is the duration (in hours).

### v4

@@ 5 Results and Discussion
$\quad$-- proxy 35.1 37.2 58.9 55.1 60.3
$\quad$-- sonar 35.4 37.7 57.9 54.7 60.9
Table 2: Average WER across different evaluation datasets. Bench: CV15.0, FLEURS and the three MGBs. IH: In-House data. Best results are shown in bold. Second best results are underlined. WER scores are reported after normalization and removing diacritics.
Scaling the Data. We scale our dataset from 100K to 500K segments and we train two models, UDW-16-16 and UDW-32-16, employing proxy-ref and sonar-sim as filtering criteria. As shown in Table 5, our UDW-32-16++ model, trained with 500K segments, performs comparably with its supervised counterpart DW-32-16++. For instance, UDW-32-16++ with proxy-ref filtering achieves a WER of 51.4%, similar to DW-32-16++’s 50.0%. Notably, our smaller student model (UDW-16-16++) trained on 500K segments results in superior performance compared to models trained under supervised filtering. For example, UDW-16-16++ with sonar-sim gives 53.9% WER compared to 56.2 for DW-16-16++. These results showcase the efficacy of our data filtering methods.
Generalization to Unseen Dialects. We evaluate our models under various unseen conditions, including five novel dialects that our models have not encountered before. Additionally, we compute category-wise  [^7] By categories we refer to dialectal and non-dialectal labels in the SADA data. error rates for both the SADA test and validation splits. We present results on the top five categories. As detailed in Table 3.1, our best model UDW-32-16++ consistently outperforms all others when averaged across these dialects.
Model NJD MTOS KHLJ HJZ UNK
Baselines
W-FT 106.4 77.0 117.8 84.2 139.9
SM4T-v1 56.5 74.0 61.9 54.6 69.6
SM4T-v2 52.6 76.2 58.2 53.3 74.7
W-M 88.2 107.9 102.8 86.7 134.0
W-L-v2 56.6 74.0 80.2 63.4 98.2
W-L-v3 53.1 73.7 70.5 61.7 89.0
DW-16-16 58.2 74.9 65.1 58.2 68.4
DW-32-16 58.7 74.9 66.7 58.7 70.1
DW-16-16++ 58.8 81.9 62.1 57.5 77.5
DW-32-16++ 51.7 67.8 57.5 52.6 67.5
@@
$\quad$-- DW-16-16 61.6 90.4 72.9 68.5 88.8
$\quad$-- DW-32-16 58.2 82.3 67.3 60.4 76.6
Ours
@@
$\quad$-- proxy 60.2 75.6 65.7 62.1 75.6
$\quad$-- sonar 64.6 59.9 84.6 72.1 66.6
@@
$\quad$-- proxy 53.4 75.7 60.0 51.4 72.8
$\quad$-- sonar 48.8 52.6 67.5 58.3 48.9
@@
$\quad$-- proxy 53.4 73.9 61.9 53.3 68.6
$\quad$-- sonar 53.7 69.4 61.4 53.7 64.6
@@
$\quad$-- proxy 48.9 66.7 60.6 48.7 66.0
$\quad$-- sonar 50.5 65.3 57.3 49.1 61.8
Table 3: WER results on top five dialects/categories on the test set of the SADA data. NJD: Najdi. MTOS: More than one speaker. KHLJ: Khaleeji. HJZ: Hijazi. UNK: Unknown. Best results are shown in bold. Second best results are underlined. WER scores are reported after normalization and removing diacritics.
Model ALG EGY JOR MAU MOR PAL UAE YEM AVG
Baselines
SM4T-v2 94.3 52.3 39.2 88.9 91.0 49.0 54.7 62.4 67.3
W-L-v2 89.9 58.1 43.2 108.3 101.1 51.9 60.8 81.4 75.6
DW-16-16 85.7 65.3 51.1 88.5 86.2 60.2 64.1 69.3 71.9
DW-32-16 86.6 65.1 51.0 89.2 87.1 59.9 63.0 69.8 72.1
No-Filter
$\quad$-- DW-32-16 93.2 66.6 48.1 95.3 86.0 58.3 63.8 68.8 73.1
Ours
UDW-16-16
$\quad$-- proxy 85.9 67.1 51.0 88.0 87.4 61.3 63.5 70.8 72.6
$\quad$-- sonar 88.4 69.8 57.3 91.3 90.9 65.7 68.4 72.8 76.2
[…]
$\quad$-- proxy 85.6 61.8 46.2 87.9 84.5 57.8 59.1 64.8 69.1
$\quad$-- sonar 82.1 58.8 45.5 86.0 85.5 53.3 57.2 63.3 67.3
Table 4: Results on the Casablanca dataset. Best results are shown in bold. Second best results are underlined. WER (↓) scores are reported after normalization and removing diacritics. We report the score on the test split of each dataset.
@@
Evaluation Dataset Baselines Ours
W-L-v2 DW-16-16 DW-32-16 UDW-16-16pr UDW-32-16pr
IID OpenBible 101.3 59.1 58.8 59.2 58.9
CommonVoice17 117.1 82.9 69.8 75.6 70.4
ALFAA 217.1 78.2 74.4 76.8 73.8
OOD DVoice 214.6 124.4 110.2 110.7 114.9
AMMI-LigAikuma 46.7 60.1 51.8 60.4 52.2
Fleurs 54.6 60.9 51.6 58.9 51.8
Table 5: WER (↓) results on the Swahili datasets. $pr$: using the proxy filtering method. Best results are shown in bold. Second best results are underlined. WER scores are reported after normalization and removing diacritics.
To further validate the effectiveness of our approach, we conduct experiments on Swahili, a low-resource language. We collect over 100 hours of labeled speech data from a variety of sources, namely OpenBible Meyer et al. (2022), CommonVoice (Swahili subset) Ardila et al. (2020), ALFAA [^8] https://github.com/besacier/ALFFA_PUBLIC/tree/master/ASR/SWAHILI, DVoice Gauthier et al. (2016), AMMI-LIGAikuma [^9] https://github.com/besacier/AMMIcourse, and FLEURS (Swahili subset) Conneau et al. (2023).
@@
We compare our distilled models to the teacher model and evaluate the performance of our unsupervised approach. The results show that our unsupervised distillation models perform on par with, or better than the supervised setup. Additionally, our distilled models outperform the teacher model by a significant margin on both familiar (IID) and novel (OOD) datasets, demonstrating the utility of our approach in extremely low-resource settings. Specifically, the UDW-32-16 model achieves a WER/CER of 58.86/14.13% on the IID OpenBible dataset, compared to the teacher model’s 101.33/44.43%. On the OOD dataset FLEURS, UDW-32-16 attains a WER/CER of 51.82/14.88, significantly outperforming the teacher model’s 54.61/14.81. Across various datasets, our distilled models consistently outperform the teacher, with UDW-32-16 showing the best results overall. Table 5 presents the WER and CER scores for the different models and datasets.
@@ 6 Conclusion
In this study, we explore methods for distilling large Whisper models into smaller, more efficient ones without relying on labeled data. Our filtering techniques bridge a gap in prior research and facilitate the creation of compact and effective speech recognition models for limited label settings. We show through a comprehensive evaluation that our models outperform both their teacher model and those using supervised distillation. Our evaluation spans a diverse range of Arabic varieties, demonstrating their generalization to linguistic diversity and their competitive performance with SOTA models twice their size. Applying our approach to Swahili datasets further validates its effectiveness for different languages. Notably, our model-based filtering methods (proxy and sonar) demonstrate superior robustness across linguistic variations. Moving forward, we aim to explore model-free approaches to further enhance the efficacy of model distillation, while including extremely low-resource languages and domains.
@@ 7 Limitations
Our main approach relies heavily on a robust reference model to serve as a proxy for filtering lower-quality pseudo labels. Specifically, we utilize SeamlessM4T-large-v2, a state-of-the-art model with 2.3 billion parameters, to generate proxy references which is then used to filter out low-quality data points. For similarity-based measures, we use SONAR Duquenne et al. (2023) to generate multimodal embeddings from speech and pseudo labels. These embeddings provide contextual similarity which is then utilized to discard low-quality pseudo labels. We use AceGPT (7B), to compute the log-likelihood of the pseudo labels which is leveraged to filter out low-quality examples.
@@
Evaluation. Arabic is a linguistically rich and complex language with over 400 million speakers Abdul-Mageed et al. (2021, 2024), resulting in its wide range of varieties and dialects. We evaluate all the models on eleven different datasets representing different varieties, including five novel dialects collected and curated by native speakers and never seen before by any models. However, our varieties do not cover all Arabic-speaking regions. We aim to address this in future work by covering more varieties and dialects.
Distillation Training Data. We distilled four variants of student models using 100K and 500K segments of which approximately 25% are filtered. We see improvement going from 100K ($\approx$100100100100 hours) to 500K ($\approx$500500500500 hours) segments. As Gandhi et al. (2023) shows, going over 1,000 hours results in a better model, we aim to study how distillation can be done under a low resource setting which is why we do not scale the data. Additionally, we also keep the WER threshold high (80) so that we remain close to a setting where no labeled data is available (even for filtering). It would be interesting, however, to see how distilled models may perform on unfiltered data in low-resource setting.
@@
Acknowledgments
We acknowledge support from Canada Research Chairs (CRC), the Natural Sciences and Engineering Research Council of Canada (NSERC; RGPIN-2018-04267), the Social Sciences and Humanities Research Council of Canada (SSHRC; 895-2020-1004; 895-2021-1008), Canadian Foundation for Innovation (CFI; 37771), Digital Research Alliance of Canada, [^10] https://alliancecan.ca and UBC Advanced Research Computing-Sockeye. [^11] https://arc.ubc.ca/ubc-arc-sockeye
@@ 2.1 SADA Dataset
Table 6: SADA stats. S is the number of segments and D is the duration (in hours). MTOS - More than one speaker.
@@ Appendix C Experiments
3.1 CER Results
Model Size CV15.0 MGB2 MGB3 MGB5 Fleurs In-house Data SADA
ALG JOR PAL UAE YEM
Baselines
Amazon -/- -/- -/- -/- -/- -/- 70.2 25.6 29.0 40.8 43.5 -/-
[…]
HuBERT 0.31 18.9 17.3 9.5 45.5 10.9 44.3 23.3 27.9 36.7 38.8 34.5
W-FT 1.5 21.9 8.1 26.9 62.3 3.4 69.6 37.2 35.4 69.1 64.8 65.7
MMS-all 1.0 80.9 13.4 34.6 45.9 6.3 78.0 55.4 75.1 78.1 76.6 38.0
[…]
SM4T-L-v1 2.3 7.3 10.5 22.6 52.1 5.1 47.8 18.8 23.1 27.4 32.5 37.8
SM4T-L-v2 2.3 3.5 8.7 18.6 53.7 4.0 52.0 14.6 17.2 23.3 30.7 41.8
[…]
W-M 0.77 13.2 18.5 39.5 88.3 5.1 69.9 21.1 24.7 52.6 52.0 74.1
W-L-v2 1.5 7.8 15.3 33.0 68.9 3.6 71.7 17.0 22.3 38.2 45.5 51.2
W-L-v3 1.5 5.2 7.6 17.3 44.6 3.2 65.4 16.3 22.7 32.7 38.9 45.6
DW-16-16 0.80 7.2 10.8 25.1 43.3 6.6 38.5 18.2 23.3 27.7 31.6 38.9
DW-32-16 1.12 5.9 8.9 21.4 40.4 4.8 33.4 14.7 19.5 22.8 28.1 49.2
[…]
DW-32-16++ 1.12 5.5 8.8 20.3 40.6 3.1 33.3 13.4 18.8 21.1 26.8 35.8
[…]
$\quad$-- DW-16-16 0.80 7.6 11.2 29.7 59.1 6.0 51.6 20.2 27.3 34.0 38.8 49.6
$\quad$-- DW-32-16 1.12 7.3 10.4 30.8 58.8 4.9 63.2 20.0 24.9 35.6 50.9 53.6
Ours
UDW-16-16 0.80
$\quad$-- nll 8.15 11.26 27.98 55.25 6.26 41.4 25.7 20.52 35.96 49.2 55.0
$\quad$-- pesq 8.41 12.11 27.69 54.88 6.89 40.34 27.41 20.16 32.55 44.17 50.1
$\quad$-- entropy 8.1 12.17 31.24 56.64 6.4 48 22.81 27.67 37.85 52.56 61.8
$\quad$-- conf 7.83 11.87 27.85 50.73 6.12 43.94 20.29 25.52 31.75 39.27 49.2
$\quad$-- proxy 7.48 11.39 26.36 49.97 7.5 42.15 23.69 19.66 30.93 41.94 46.2
$\quad$-- sonar 8.04 11.86 28.66 49.21 7.06 43.48 22.61 27.43 32.89 36.13 45.6
[…]
$\quad$-- nll 6.24 10.12 25.39 55.53 4.47 35.85 20.88 16.04 30.49 41.38 46.1
$\quad$-- pesq 7.5 10.6 26.4 51.0 5.3 43.2 17.2 22.7 30.9 36.4 48.1
$\quad$-- entropy 6.53 10.34 28.71 66.87 4.34 84.02 21.07 31.08 44.23 54.25 52.3
$\quad$-- conf 6.46 9.79 23.42 48.61 4.73 36.08 21.99 25.82 17.7 42.82 41.3
$\quad$-- proxy 6.17 9.87 22.45 46.05 6.23 36.11 15.6 20.88 25.69 29.49 41.9
$\quad$-- sonar 5.62 8.98 23.4 46.97 4.41 37 15.11 19.56 24.29 28.24 35.6
[…]
$\quad$-- proxy 4.8 10.4 24.3 48.0 4.8 41.6 16.3 21.1 27.5 33.0 35.3
$\quad$-- sonar 6.1 9.8 24.2 47.1 5.0 38.9 17.2 21.4 26.2 29.5 33.96
UDW-32-16-++ 1.12
$\quad$-- proxy 5.8 9.2 21.1 44.1 4.2 37.1 14.5 20.1 23.1 29.2 31.4
$\quad$-- sonar 5.3 9.9 22.4 44.2 3.9 34.6 19.2 15.1 23.2 27.0 33.4
Table 7: CER ($\downarrow$) scores after normalization and removing diacritics. All baseline distilled models (DW-) are trained with a filtering threshold of 80 if not specified. Best results are shown in bold. Second best results are underlined. We report the score on the test split of each dataset. Abbreviations. W - Whisper, FT - Finetuned, M - Medium, L - Large, S - Small, U - Unsupervised, D - Distil, nll - negative log likelihood, conf - confidence score.
Model Split NJD MTOS KHLJ HJZ UNK
Baslines
W-FT Test 77.5 51.8 85.4 61.5 112.2
Valid 52.6 41.1 100.3 89.7 107.6
SM4T-v1 Test 30.9 46.0 32.2 29.0 39.4
Valid 28.1 44.0 31.5 30.9 35.2
SM4T-v2 Test 31.1 53.1 30.4 32.0 45.1
Valid 30.7 53.7 35.3 30.3 34.4
W-M Test 65.8 79.3 77.0 59.7 122.2
Valid 56.9 75.1 62.9 52.0 106.5
W-L-v2 Test 39.9 57.4 54.4 39.6 80.7
Valid 41.4 55.4 44.9 43.6 67.1
W-L-v3 Test 31.6 53.7 44.1 38.6 61.3
Valid 30.2 47.7 39.2 27.2 49.2
DW-16-16 Test 30.8 47.6 32.7 30.7 39.8
Valid 31.4 44.7 35.2 32.8 39.8
DW-32-16 Test 35.8 60.1 38.7 34.1 44.5
Valid 34.8 54.1 37.6 38.2 40.1
DW-16-16++ Test 30.9 50.7 31.5 31.0 46.8
Valid 29.8 43.8 31.8 33.0 41.0
DW-32-16++ Test 28.3 43.1 29.4 28.6 41.3
Valid 27.3 38.3 34.5 28.1 43.0
[…]
$\quad$-- DW-16-16 Test 34.8 59.7 41.4 42.3 63.0
Valid 38.9 53.4 41.9 37.7 54.3
$\quad$-- DW-32-16 Test 42.8 63.9 47.0 45.9 63.2
Valid 35.2 54.9 43.3 36.5 49.6
Ours
UDW-16-16
$\quad$-- proxy Test 35.5 55.6 38.9 39.6 52.0
Valid 34.0 50.9 39.1 37.1 41.2
$\quad$-- sonar Test 35.8 30.3 55.7 38.8 36.8
Valid 35.9 39.3 52.4 38.7 36.7
[…]
$\quad$-- proxy Test 31.1 54.0 32.1 30.8 46.0
Valid 29.3 44.3 29.1 28.6 36.6
$\quad$-- sonar Test 25.4 23.6 45.9 29.9 25.5
Valid 26.0 26.7 44.1 30.3 29.5
[…]
$\quad$-- proxy Test 29.7 48.8 33.8 29.6 42.8
Valid 27.8 42.0 34.3 32.2 41.7
$\quad$-- sonar Test 28.4 43.3 30.8 27.5 37.0
Valid 27.5 40.3 32.4 30.7 35.8
UDW-32-16++
[…]
Valid 25.3 37.4 30.1 25.6 37.4
$\quad$-- sonar Test 26.3 40.8 28.3 24.8 34.5
Valid 25.3 37.0 30.2 29.8 34.2
Table 8: CER ($\downarrow$) results on top five dialects/categories in SADA data. Best results are shown in bold. Second best results are underlined. The scores are reported after normalization and removing diacritics.
Model NJD MTOS KHLJ HJZ UNK
Baselines
W-FT 77.1 63.4 139.4 119.1 140.3
SM4T-v1 51.9 68.7 61.7 54.2 62.3
SM4T-v2 52.2 75.8 65.1 51.1 59.8
W-M 80.4 102.8 89.5 72.9 127.7
W-L-v2 60.9 72.9 67.7 64.5 68.0
W-L-v3 49.3 65.5 67.5 46.5 67.7
DW-16-16 59.4 70.6 66.2 61.1 69.9
DW-32-16 58.3 69.7 67.5 62.7 68.3
DW-16-16++ 56.8 72.0 62.3 60.2 75.2
DW-32-16++ 50.3 61.8 62.3 53.7 66.4
[…]
$\quad$-- DW-16-16 64.8 80.3 71.4 65.5 77.0
$\quad$-- DW-32-16 57.9 73.7 68.6 56.8 72.3
Ours
UDW-16-16
$\quad$-- proxy 59.3 70.7 66.5 61.4 68.7
$\quad$-- sonar 64.8 67.5 78.1 69.9 65.3
[…]
$\quad$-- proxy 51.0 65.9 58.1 53.7 64.9
$\quad$-- sonar 49.2 51.6 62.5 58.9 52.6
[…]
$\quad$-- proxy 52.7 66.6 62.0 55.7 67.3
$\quad$-- sonar 53.6 64.4 62.3 55.7 63.7
[…]
$\quad$-- proxy 49.0 60.4 57.6 50.7 60.7
$\quad$-- sonar 49.3 58.8 58.1 53.6 61.1
Table 9: WER ($\downarrow$) results on top five dialects/categories on the validation set of the SADA data. Best results are shown in bold. Second best results are underlined. WER scores are reported after normalization and removing diacritics.
Model Bench SADA2022 IH
Test Valid Test Valid
Baselines
HuBERT 20.4 22.7 34.5 31.9 34.2
W-FT 24.5 28.6 65.7 56.2 55.2
[…]
SM4T-v2 17.7 19.4 41.8 40.8 27.6
W-M 32.9 38.0 74.1 66.7 44.1
W-L-v2 25.7 29.7 51.2 48.5 38.9
W-L-v3 15.6 17.1 45.6 39.3 35.2
DW-16-16 18.6 20.5 38.9 37.8 27.8
DW-32-16 16.3 17.9 47.3 43.9 23.7
DW-16-16++ 17.8 19.4 40.6 36.6 27.2
DW-32-16++ 15.7 17.1 35.8 33.4 22.7
[…]
$\quad$-- DW-16-16 22.7 25.1 49.6 45.9 34.4
$\quad$-- DW-32-16 22.5 25.8 53.6 45.0 38.9
Ours
UDW-16-16
$\quad$-- proxy 20.5 22.1 46.2 42.2 31.7
$\quad$-- sonar 21.0 22.8 45.6 43.5 32.5
[…]
$\quad$-- proxy 18.2 19.6 41.9 36.1 25.6
$\quad$-- sonar 17.9 19.7 35.6 34.7 24.8
[…]
$\quad$-- proxy 18.4 21.2 39.2 35.3 27.9
$\quad$-- sonar 18.4 19.9 35.4 34.0 26.6
[…]
$\quad$-- proxy 17.1 18.7 33.6 31.4 23.8
$\quad$-- sonar 16.9 18.5 33.1 31.3 24.8
Table 10: Average CER ($\downarrow$) across different evaluation datasets. Bench: CV15.0, FLEURS and the three MGBs. Best results are shown in bold. Second best results are underlined. The scores are reported after normalization and removing diacritics.
Evaluation Dataset Baselines Ours
W-L-v2 DW-16-16 DW-32-16 UDW-16-16pr UDW-32-16pr
IID OpenBible 44.4 14.0 13.8 14.0 14.1
CommonVoice17 60.1 35.0 24.8 29.2 25.4
ALFAA 143.2 28.2 25.7 27.2 26.5
OOD DVoice 144.6 74.1 62.6 62.4 69.1
AMMI-LigAikuma 13.0 18.0 14.4 18.5 14.4
Fleurs 14.8 18.9 14.8 18.5 14.9
Table 11: CER (↓) results on the Swahili datasets. $pr$: using the proxy filtering method. Best results are shown in bold. Second best results are underlined. WER scores are reported after normalization and removing diacritics
Model ALG EGY JOR MAU MOR PAL UAE YEM AVG
Baselines
SM4T-v2 53.48 26.12 13.15 52.20 54.96 18.20 22.71 27.07 34.44
W-L-v2 58.63 30.28 20.37 79.66 63.21 25.70 38.06 51.49 46.83
DW-16-16 40.08 31.80 19.11 49.83 42.16 24.10 26.99 30.53 33.64
DW-32-16 44.45 32.80 19.27 49.95 43.46 26.43 26.26 34.03 35.12
No-Filter
$\quad$-- DW-32-16 61.50 43.52 18.41 64.19 51.36 29.44 36.97 41.75 43.95
Ours
UDW-16-16
$\quad$-- proxy 48.30 39.79 20.21 53.06 45.92 25.69 29.15 37.13 38.01
$\quad$-- sonar 43.94 36.24 23.60 55.10 50.14 28.77 31.05 34.65 38.63
[…]
$\quad$-- proxy 40.72 29.89 16.23 47.03 41.45 23.42 23.72 27.26 31.80
$\quad$-- sonar 38.34 28.61 16.02 50.02 44.94 19.87 23.13 27.17 31.79
Table 12: CER results on the Casablanca dataset. The best results are shown in bold. The second-best results are underlined. CER ($\downarrow$) scores are reported after normalization and removing diacritics. We report the score on the test split of each dataset.